OpenStudy (anonymous):

what is the square root of 36 and 180

OpenStudy (anonymous):

\[6\times6=36 \] \[\sqrt{36}=6\] \[\sqrt{180}=13.416\]

OpenStudy (anonymous):

\[\sqrt{180}=6\sqrt{5}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

YW